Find three consecutive odd integers such that the product of three and the third integer is twenty-one more than two times the sum of the first two numbers.
Thank you, Thank you

Answer by Paul(988)

First = x
second = x+2
third = x+4
3(x+4)=21+2(x+x+2)
3x+12=21+4x+4
3x+12=25+4x
x=-13
-13+2=-11
-13+4=-9
Hence, the three consective numbers are -13, -11, and -9